A concentric tube heat exchanger of length L = 2 m
Answer #1
Concepts and reason

To solve the problem, relation between dimensionless numbers is to be understood. Also, the concept of heat exchanger analysis using effectiveness-number of transfer unit’s methods is also to be understood.

Convection:

Whenever the heat exchange is due to the movement of fluid particles it is called as convective heat transfer. The fluid particles transmit heat to a location at a lower temperature from a location at a higher temperature. Convection heat transfer requires a medium to propagate heat exchange.

Dimensionless numbers:

Dimensionless numbers are the numbers which do not have any physical dimension assigned to them.

Heat exchanger:

Heat exchangers are devices used for transfer of heat energy from one object to another. There are types of heat exchangers:

1.Parallel-flow heat exchanger.

2.Counter-flow heat exchanger.

Parallel-flow heat exchanger:

This is such a heat exchange in which the flow direction of the both the hot and cold fluids that flows inside the tube is unidirectional.

Counter-flow heat exchanger:

This is such a heat exchange in which the flow direction of the one of the fluid that flows inside the tube is opposite to the respective cold fluid.

To obtain solution for the given problem, initially calculate the Reynolds number for the cold fluid and determine the type of flow. Depending on the type of flow use the Nusselt number expression to calculate the convective heat transfer coefficient of the cold fluid. Similarly calculate the Reynolds number for the hot fluid and determine the type of flow.

Use the Nusselt number corresponding to the flow and determine the value of convective heat transfer coefficient of the hot fluid. Use the overall heat transfer coefficient expression, substitute the corresponding values and obtain the value of overall heat transfer coefficient.

Calculate the mass flow rate of cold and hot fluid. Further calculate the heat capacity of each fluid. Also determine the heat capacity ratio and NTU
values. Use the expression for effectiveness expression for counter flow heat exchanger substitute the respective values and obtain the value of effectiveness of the heat exchanger.

Later use the expression for effectiveness in terms of temperatures, substitute the values and calculate the mean outlet temperature of the cold fluid. Similarly repeat the effectiveness -NTU
procedure for parallel flow heat exchanger.

Fundamentals

The expression for liner interpolation to determine the value of an unknown data say, yy , at a given value of xx is given as

y=(xx1)(y2y1)(x2x1)+y1y = (x - {x_1})\frac{{\left( {{y_2} - {y_1}} \right)}}{{\left( {{x_2} - {x_1}} \right)}} + {y_1}

Here, (x1,y1)\left( {{x_1},{y_1}} \right) and (x2,y2)\left( {{x_2},{y_2}} \right) are the nearest data points taken from the given set of data.

Hydraulic diameter:

It is the effective diameter used for fluid flow in a pipe, conduit or duct of any shape. Mathematically it is given as,

4Ag
e

Here, is the hydraulic diameter,A
is the cross-section area of the duct or pipe and Р
is the wetted perimeter of the duct or pipe.

Hydraulic diameter of the annular space:

For a concentric tube with an outer diameter and inner diameter , the hydraulic diameter is calculated as:

4 (T/4)(D-D)
(D-D)
- D-D

Here, is the hydraulic diameter, is the outer diameter of the concentric tube and is the inner diameter of the concentric tube.

Reynolds Number:

For flow through pipe, Reynolds number is given by,

Rep

Here, is the Reynolds number for flow through pipe, is the mean fluid velocity, is the diameter of the pipe, is the density of the fluid and is the coefficient of dynamic viscosity of the fluid.

It is also expressed as,

uD
Rep
V

Here, is the kinematic viscosity.

The critical Reynolds number corresponding to the onset of turbulence is given as:

Re
cri2300

For Reynolds number below 2300
the flow is considered to be laminar and above the critical Reynolds number the flow is considered as turbulent.

Nusselt number:

Nusselt number provides the measure of convective heat transfer occurring at the surface. Mathematical expression for Nusselt number for flow through pipe is given as,

hD
Nup

Here,Nup
is the Nusselt number for the flow through pipe,h
is the convective heat transfer coefficient and k
is the thermal conductivity of the fluid.

Nusselt number varies with the type of the flow.

Nusselt number for fully developed laminar flow in a circular tube annulus with one surface insulated and the other at constant temperature and with the ratio of inner diameter to outer diameter of pipe as:

Nu, 5.74

Here,Nu,
is the Nusselt number of the inner surface of the pipe.

For turbulent flow through pipe the Nusselt number is a function of Reynolds number and Prandtl number and is given as,

Nup 0.023(Re) (Pr)4

Here,Nup
is the Nusselt number for flow through pipe and Pr
is the Prandtl number.

Overall heat transfer coefficient:

Overall heat transfer coefficient is related to total thermal resistance of heat transfer between two fluids. For tubular heat exchangers it is given by,

R In(D D)R
hA А
1
+
UA
А, hА
2лk!

Here, is the overall hear transfer coefficient, is the heat transfer coefficient of the inner fluid and is the heat transfer coefficient of the outer fluid, is the length of the heat exchanger, is the fouling resistance of the inner surface of heat exchanger, R
is the fouling resistance of the outer surface of heat exchanger, A
is the surface area, is the area of inner surface of heat exchanger and A,
is the area of outer surface of heat exchanger.

Neglecting the resistance due to fouling factors and conduction resistance, simplified form of overall heat transfer coefficient is given by,

1
U
1
h

Mass flow rate:

For flow through pipe, mass flow rate for a fluid is given as,

|m = pAu

Here, m
is the mass flow rate of the fluid, is the density of the fluid, A
is the cross-sectional area of the pipe and и
is the mean velocity of the pipe.

Heat Capacity ratio:

It is defined as the ratio of minimum heat capacity of the fluid to the maximum heat capacity of the fluid and is given as:

С.
R
С.
min
max

Here, R
is the heat capacity ratio, С.
min
is the minimum heat capacity of the fluid and С.
max
is the maximum heat capacity of the fluid.

Number of transfer units:

Number of transfer units is a dimensionless parameter used for heat exchanger analysis and is given by,

NTU=UA
CA
min

Here, NTU
is the number of transfer units.

Effectiveness of the heat exchanger:

Effectiveness of a heat exchanger is defined as the ratio of actual heat transfer rate of a heat exchanger to the maximum possible heat transfer rate. It is given as,

пах

Here, is the effectiveness of the heat exchanger, is the actual heat transfer rate of the heat transfer rate and max
пах
is the maximum possible heat transfer rate.

It is also expressed in terms of temperatures as:

Or,

Here,со
is the outlet temperature of cold fluid, ci
is the inlet temperature of cold fluid, T
ho
is the outlet temperature of hot fluid, is the inlet temperature of hot fluid, is the heat capacity of cold fluid, is the heat capacity of hot fluid and С.
max
is maximum heat capacity of either hot or cold fluid.

(a)

Write the expression of temperature conversion from Celsius to kelvin.

T=T(C)+273T = T\left( {^\circ {\rm{C}}} \right) + 273

Here, TT is the temperature in kelvin and T(C)T\left( {^\circ {\rm{C}}} \right) is the temperature in degree Celsius.

Substitute 50C50^\circ {\rm{C}} for T(C)T\left( {^\circ {\rm{C}}} \right) .

T=50C+273=323K\begin{array}{c}\\T = 50^\circ {\rm{C}} + 273\\\\ = 323{\rm{K}}\\\end{array}

Understand that the value of properties of water given in the table are at 320K320{\rm{K}} and 325K325{\rm{K}} . To calculate the values at 323K323{\rm{K}} use the method of interpolation.

Consider Tx=325K{T_x} = 325{\rm{K}} and Ty=320K{T_y} = 320{\rm{K}} , and obtain the corresponding values of coefficient of dynamic viscosity and specific heat of water from the table of Thermophysical Properties of Saturated Water.

μx=577×106Ns/m2μy=528×106Ns/m2cpx=4.180kJ/kgKcpy=4.182kJ/kgK\begin{array}{l}\\{\mu _x} = {\rm{577}} \times {10^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2}\\\\{\mu _y} = {\rm{528}} \times {10^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2}\\\\{c_{{p_x}}} = {\rm{4}}{\rm{.180 kJ/kg}} \cdot {\rm{K}}\\\\{{\rm{c}}_{{p_y}}} = {\rm{4}}{\rm{.182 kJ/kg}} \cdot {\rm{K}}\\\end{array}

Here, μx{\mu _x} is the coefficient of dynamic viscosity at 320K320{\rm{K}} , μy{\mu _y} is the coefficient of dynamic viscosity at 325K325{\rm{K}} , cpx{c_{{p_x}}} is the specific heat of water at 320K320{\rm{K}} and cpy{c_{{p_y}}} is the specific heat of water at 325K325{\rm{K}} .

Further obtain the thermal conductivity and Prandtl Number corresponding to Tx{T_x} and Ty{T_y} from the table of Thermophysical Properties of Saturated Water.

kx=640×103W/mKky=645×103W/mKPrx=3.77Pry=3.42\begin{array}{l}\\{k_x} = {\rm{640}} \times {10^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}}\\\\{k_y} = {\rm{645}} \times {10^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}}\\\\{\rm{P}}{{\rm{r}}_x} = {\rm{3}}{\rm{.77}}\\\\{\rm{P}}{{\rm{r}}_y} = {\rm{3}}{\rm{.42}}\\\end{array}

Here, kx{k_x} is the thermal conductivity at 320K320{\rm{K}} , ky{k_y} is the thermal conductivity at 325K325{\rm{K}} , Prx{\rm{P}}{{\rm{r}}_x} is the Prandtl Number at 320K320{\rm{K}} and Pry{\rm{P}}{{\rm{r}}_y} is the Prandtl Number at 325K325{\rm{K}} .

Write the expression of interpolation for coefficient of dynamic viscosity of water at 323K323{\rm{K}} .

μh=(TTx)(μyμx)(TyTx)+μx{\mu _h} = (T - {T_x})\frac{{\left( {{\mu _y} - {\mu _x}} \right)}}{{\left( {{T_y} - {T_x}} \right)}} + {\mu _x}

Here, is the dynamic viscosity of the hot fluid.

Substitute 323K323{\rm{K}} for TT , 577×106Ns/m2{\rm{577}} \times {10^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2} for μx{\mu _x} , 528×106Ns/m2{\rm{528}} \times {10^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2} for μy{\mu _y} , 325K325{\rm{K}} for Ty{T_y} and 320K320{\rm{K}} for Tx{T_x} .

μh=(323K320K)(528×106Ns/m2577×106Ns/m2)(325K320K)+577×106Ns/m2=(3)(49×1065)+577×106=29.4×106+577×106=547.6×106Ns/m2\begin{array}{c}\\\;{\mu _h} = (323{\rm{K}} - 320{\rm{K}})\frac{{\left( {{\rm{528}} \times {{10}^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2} - {\rm{577}} \times {{10}^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2}} \right)}}{{\left( {325{\rm{K}} - 320{\rm{K}}} \right)}} + {\rm{577}} \times {10^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2}\\\\ = \left( 3 \right)\left( {\frac{{ - 49 \times {{10}^{ - 6}}}}{5}} \right) + {\rm{577}} \times {10^{ - 6}}\\\\ = - 29.4 \times {10^{ - 6}} + {\rm{577}} \times {10^{ - 6}}\\\\ = 547.6 \times {10^{ - 6}}{\rm{ N}} \cdot {\rm{s/}}{{\rm{m}}^2}\\\end{array}

Write the expression of interpolation for specific heat of water of water at 323K323{\rm{K}} .

cp,h=(TTx)(cpycpx)(TyTx)+cpx{c_{p,h}} = (T - {T_x})\frac{{\left( {{c_{{p_y}}} - {c_{{p_x}}}} \right)}}{{\left( {{T_y} - {T_x}} \right)}} + {c_{{p_x}}}

Here, С ph
is the specific heat of the hot fluid.

Substitute 323K323{\rm{K}} for TT , 4.180kJ/kgK{\rm{4}}{\rm{.180 kJ/kg}} \cdot {\rm{K}} for cpx{c_{{p_x}}} , 4.182kJ/kgK{\rm{4}}{\rm{.182 kJ/kg}} \cdot {\rm{K}} for cpy{{\rm{c}}_{{p_y}}} , 325K325{\rm{K}} for Ty{T_y} and 320K320{\rm{K}} for Tx{T_x} .

cp,h=(323K320K)(4.182kJ/kgK4.180kJ/kgK)(325K320K)+4.180kJ/kgK=(3)(0.0025)+4.180=0.0012+4.180=4.181kJ/kgK\begin{array}{c}\\\;{c_{p,h}} = (323{\rm{K}} - 320{\rm{K}})\frac{{\left( {{\rm{4}}{\rm{.182 kJ/kg}} \cdot {\rm{K}} - {\rm{4}}{\rm{.180 kJ/kg}} \cdot {\rm{K}}} \right)}}{{\left( {325{\rm{K}} - 320{\rm{K}}} \right)}} + {\rm{4}}{\rm{.180 kJ/kg}} \cdot {\rm{K}}\\\\ = \left( 3 \right)\left( {\frac{{0.002}}{5}} \right) + {\rm{4}}{\rm{.180}}\\\\ = 0.0012 + {\rm{4}}{\rm{.180}}\\\\ = {\rm{4}}{\rm{.181 kJ/kg}} \cdot {\rm{K}}\\\end{array}

Write the expression of interpolation for thermal conductivity of water at 323K323{\rm{K}} .

(kh)o=(TTx)(kykx)(TyTx)+kx{\left( {{k_h}} \right)_o} = (T - {T_x})\frac{{\left( {{k_y} - {k_x}} \right)}}{{\left( {{T_y} - {T_x}} \right)}} + {k_x}

Here, (kh)
is the thermal conductivity of hot fluid flowing in annular region of the tube.

Substitute 323K323{\rm{K}} for TT , 640×103W/mK{\rm{640}} \times {10^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}} for kx{k_x} , 645×103W/mK{\rm{645}} \times {10^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}} for ky{k_y} , 325K325{\rm{K}} for Ty{T_y} and 320K320{\rm{K}} for Tx{T_x} .

(kh)o=(323K320K)(645×103W/mK640×103W/mK)(325K320K)+640×103W/mK=(3)(5×1035)+640×103=(3×103)+(640×103)=643×103W/mK\begin{array}{c}\\\;{\left( {{k_h}} \right)_o} = (323{\rm{K}} - 320{\rm{K}})\frac{{\left( {{\rm{645}} \times {{10}^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}} - {\rm{640}} \times {{10}^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}}} \right)}}{{\left( {325{\rm{K}} - 320{\rm{K}}} \right)}} + {\rm{640}} \times {10^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}}\\\\ = \left( 3 \right)\left( {\frac{{5 \times {{10}^{ - 3}}}}{5}} \right) + {\rm{640}} \times {10^{ - 3}}\\\\ = \left( {3{\kern 1pt} \times {{10}^{ - 3}}} \right) + \left( {{\rm{640}} \times {{10}^{ - 3}}} \right)\\\\ = {\rm{643}} \times {10^{ - 3}}{\rm{ W/m}} \cdot {\rm{K}}\\\end{array}

Write the expression of interpolation for Prandtl Number at 323K323{\rm{K}} .

Pr=(TTx)(PryPrx)(TyTx)+Prx{\rm{Pr}} = (T - {T_x})\frac{{\left( {{\rm{P}}{{\rm{r}}_y} - {\rm{P}}{{\rm{r}}_x}} \right)}}{{\left( {{T_y} - {T_x}} \right)}} + {\rm{P}}{{\rm{r}}_x}

Substitute 323K323{\rm{K}} for TT , 3.77{\rm{3}}{\rm{.77}} for Prx{\rm{P}}{{\rm{r}}_x} , 3.42{\rm{3}}{\rm{.42}} for Pry{\rm{P}}{{\rm{r}}_y} , 325K325{\rm{K}} for Ty{T_y} and 320K320{\rm{K}} for Tx{T_x} .

Pr=(323K320K)(3.423.77)(325K320K)+3.77=(3)(0.355)+3.77=(0.21)+3.77=3.56\begin{array}{c}\\\;{\rm{Pr}} = (323{\rm{K}} - 320{\rm{K}})\frac{{\left( {{\rm{3}}{\rm{.42}} - {\rm{3}}{\rm{.77}}} \right)}}{{\left( {325{\rm{K}} - 320{\rm{K}}} \right)}} + {\rm{3}}{\rm{.77}}\\\\ = \left( 3 \right)\left( {\frac{{ - 0.35}}{5}} \right) + {\rm{3}}{\rm{.77}}\\\\ = \left( { - 0.21} \right) + {\rm{3}}{\rm{.77}}\\\\ = 3.56\\\end{array}

Obtain the density of water at 323K323{\rm{K}} from its property table.

ρ=988.1kg/m3\rho = 988.1{\rm{ kg/}}{{\rm{m}}^3}

Write the expression for Reynolds number for flow through pipe.

и.
|(Re2) D

Here, is the mean velocity of the cold fluid, is the inner diameter of the tube, is the kinematic viscosity of the cold fluid and (Re,)
is the Reynolds number of the cold fluid.

Substitute for и
m.e
, for and for .

(0.1 m/s) 10 mm0.001 m
(Re)
1 mm
10x10 m2/s
(103 m2/s)
(103 m2/s)
=100

Understand that, the value (Re,)
is less than 2300
. Hence, the nature of flow through the inner tube is laminar.

Write the expression of Nusselt number for inner pipe.

Nui=hiDi(kc)iN{u_i} = \frac{{{h_i}{D_i}}}{{{{\left( {{k_c}} \right)}_i}}}

Here, is the heat transfer coefficient for fluid in the inner pipe of concentric tube, is the diameter of the inner pipe, is the thermal conductivity of cold fluid flowing in inner pipe and Nu,
is the Nusselt number of the inner surface of the pipe.

Write the expression of ratio of inner diameter to outer diameter of pipe.

n=DiDon = \frac{{{D_i}}}{{{D_o}}}

Here, nn is the ratio of inner diameter to outer diameter of pipe.

Substitute 10mm10{\rm{ mm}} for Di{D_i} and 20mm20{\rm{ mm}} for Do{D_o} .

n=10mm20mm=0.5\begin{array}{c}\\n = \frac{{10{\rm{ mm}}}}{{20{\rm{ mm}}}}\\\\ = 0.5\\\end{array}

Write the value for Nusselt number for fully developed laminar flow in a circular tube annulus with one surface insulated and the other at constant temperature and with the ratio of inner diameter to outer diameter as .

Substitute h,D,
(k.)
for Nu,
.

Substitute for and 0.25 W/mK
for (k)
.

hi(10mm(0.001m1mm))(0.25W/mK)=5.74\frac{{{h_i}\left( {{\rm{10 mm}}\left( {\frac{{0.001{\rm{ m}}}}{{1{\rm{ mm}}}}} \right)} \right)}}{{\left( {0.25{\rm{ W/mK}}} \right)}} = 5.74

Solve for hi{h_i} .

(hi)(0.01)=1.435hi=143.5W/m2K\begin{array}{l}\\\left( {{h_i}} \right)\left( {0.01} \right) = 1.435\\\\{h_i} = 143.5{\rm{ W/}}{{\rm{m}}^2}K\\\end{array}

Write the expression for hydraulic diameter for the annular space.

D D-D

Substitute for and for .

Write the expression for Reynolds number for flow through annular region of the pipe.

Here, u
mh
is the mean velocity of the hot fluid, is the hydraulic diameter of the tube, is the density of the hot fluid and (Re,)
is the Reynolds number of the hot fluid.

Substitute for , for , for and for .

As the Reynolds number is greater than 2300, the flow is considered to be turbulent.

Write the expression of Nusselt number for annular region of concentric tube.

Nuo=hoDo(kh)oN{u_o} = \frac{{{h_o}{D_o}}}{{{{\left( {{k_h}} \right)}_o}}}

Here, NuoN{u_o} is the Nusselt number for annular region of concentric tube, is the heat transfer coefficient for fluid in the annular region of concentric tube and is the hydraulic diameter of the annular region.

Write the expression for Nusselt number for turbulent flow.

Substitute for .

Substitute for , for Pr
, for and for .

Write the expression for overall heat transfer coefficient.

Substitute for and for .

U=11(143.5W/m2K)+1(1723.2W/m2K)=16.96864115×103+5.803156917×104=17.54895×103=132.5W/m2K\begin{array}{c}\\U = \frac{1}{{\frac{1}{{\left( {143.5{\rm{ W/}}{{\rm{m}}^2}{\rm{K}}} \right)}} + \frac{1}{{\left( {1723.2{\rm{ W/}}{{\rm{m}}^2}{\rm{K}}} \right)}}}}\\\\ = \frac{1}{{6.96864115 \times {{10}^{ - 3}} + 5.803156917 \times {{10}^{ - 4}}}}\\\\ = \frac{1}{{7.54895 \times {{10}^{ - 3}}}}\\\\ = 132.5{\rm{ W/}}{{\rm{m}}^2}{\rm{K}}\\\end{array}

(b)

Write the expression to calculate the area of the inner tube.

Here, is the cross-sectional area and is the inner diameter of the tube.

Substitute for .

Write the expression to calculate the area of the annulus of tube.

Here, is the cross-sectional area is the outer diameter of the tube and is the inner diameter of the tube.

Substitute for and for .

Write the expression for mass flow rate of cold fluid, that is pharmaceutical product.

Here, is the density of the cold fluid, is cross sectional area of the inner tube, is the mean velocity of the cold fluid and is the density of the cold fluid.

Substitute for ,for and for .

Write the expression for mass flow rate of hot fluid, that is water.

Here, is the density of the hot fluid, is cross sectional area of the annulus of tube, is the mean velocity of the hot fluid and is the density of the hot fluid.

Substitute for ,for and for u
mh
.

, (987.9 kg/m)(235.62x10 m2)(0.2 m/s)
=0.0466 kg/s

Write the expression for heat capacity of the cold fluid.

Here, is the heat capacity of the cold fluid, is the mass flow rate of the cold fluid and Сре
is the specific heat of the cold fluid.

Substitute 0.00864 kg/s
for and for .

Write the expression for heat capacity of the hot fluid.

Here,С,
h
is the heat capacity of the hot fluid and is the mass flow rate of the hot fluid.

Substitute for and 4.181kJ/kgK4.181{\rm{ kJ/kgK}} for .

Ch=(0.0466kg/s)(4.181kJ/kgK(1000J1kJ))=194.83W/K\begin{array}{c}\\{C_h} = \left( {0.0466{\rm{ kg/s}}} \right)\left( {4.181{\rm{ kJ/kgK}}\left( {\frac{{1000{\rm{ J}}}}{{1{\rm{ kJ}}}}} \right)} \right)\\\\ = 194.83{\rm{ W/K}}\\\end{array}

Write the expression for heat capacity ratio.

Substitute for and for .

R=CcChR = \frac{{{C_c}}}{{{C_h}}}

Substitute for and for С,
h
.

Write the expression of surface area of the tube.

A=πDiLA = \pi {D_i}L

Here, is the length of the tube.

Write the expression of .

NTU=UACminNTU = \frac{{UA}}{{{C_{\min }}}}

Substitute for and πDiL\pi {D_i}L for AA .

NTU=U(πDiL)CcNTU = \frac{{U\left( {\pi {D_i}L} \right)}}{{{C_c}}}

Substitute for ,for ,for and for .

Write the expression for effectiveness in terms of NTU
for counter flow heat exchanger.

Substitute for R
and 0.392
for .

Write the expression for effectiveness in terms of temperature.

Substitute for .

Substitute for ci
, for and for .

0.32=(Tco)(293K)(333K)(293K)Tco(293K)=12.8KTco=305.8K\begin{array}{l}\\0.32 = \frac{{\left( {{T_{co}}} \right) - \left( {293{\rm{ K}}} \right)}}{{\left( {333{\rm{ K}}} \right) - \left( {293{\rm{ K}}} \right)}}\\\\{T_{co}} - \left( {293{\rm{ K}}} \right) = 12.8{\rm{ K}}\\\\{T_{co}} = 305.8{\rm{ K}}\\\end{array}

Convert the unit of temperature from Kelvin to Celsius.

Tco=305.8K=(305.8273)C=32.8C\begin{array}{c}\\{T_{co}} = 305.8{\rm{ K}}\\\\{\rm{ = }}\left( {305.8 - 273} \right)^\circ {\rm{C}}\\\\ = 32.{\rm{8}}^\circ {\rm{C}}\\\end{array}

(c)

Write the expression for effectiveness in terms of for parallel flow heat exchanger.

Substitute for and for .

1-cxp-(0.392)(1+0.109)]
= 3
1+0.109
0.353
1.109
=0.318

Write the expression for effectiveness in terms of temperature.

Substitute for .

Substitute for , for and for .

0.318=(Tco)(293K)(333K)(293K)Tco(293K)=12.72KTco=305.72K\begin{array}{l}\\0.318 = \frac{{\left( {{T_{co}}} \right) - \left( {293{\rm{ K}}} \right)}}{{\left( {333{\rm{ K}}} \right) - \left( {293{\rm{ K}}} \right)}}\\\\{T_{co}} - \left( {293{\rm{ K}}} \right) = 12.72{\rm{ K}}\\\\{T_{co}} = 305.72{\rm{ K}}\\\end{array}

Convert this temperature from Kelvin to degree Celsius.

Tco=305.72K=(305.72273)C=32.72C\begin{array}{c}\\{T_{co}} = 305.72{\rm{ K}}\\\\{\rm{ = }}\left( {305.72 - 273} \right)^\circ {\rm{C}}\\\\ = 32.72^\circ {\rm{C}}\\\end{array}

Ans: Part a

The value of overall heat transfer coefficient is .

Part b

The mean outlet temperature of pharmaceutical product when operating in counter flow mode is .

Part c

The mean outlet temperature of pharmaceutical product when operating in parallel flow mode is 32.72° C
.

answered by: ping